### v1

# OBD-Finder: Explainable Coarse-to-Fine Text-Centric Oracle Bone Duplicates Discovery

Abstract
Oracle Bone Inscription (OBI) is the earliest systematic writing system in China, while the identification of Oracle Bone (OB) duplicates is a fundamental issue in OBI research. In this work, we design a progressive OB duplicate discovery framework that combines unsupervised low-level keypoints matching with high-level text-centric content-based matching to refine and rank the candidate OB duplicates with semantic awareness and interpretability. We compare our approach with state-of-the-art content-based image retrieval and image matching methods, showing that our approach yields comparable recall performance and the highest simplified mean reciprocal rank scores for both Top-5 and Top-15 retrieval results, and with significantly accelerated computation efficiency. We have discovered over 60 pairs of new OB duplicates in real-world deployment, which were missed by OBI researchers for decades. The models, video illustration and demonstration of this work are available at: https://github.com/cszhangLMU/OBD-Finder/.
Keywords:
Oracle Bone duplicates Image matching Image retrieval.
1 Introduction
Oracle Bone Inscription (OBIs) was used in the late Shang Dynasty more than 3000 years ago for divination and recording purposes. But from then on, these Oracle Bones (OBs) had been buried underground for thousands of years, until they were rediscovered in 1899 for containing inscribed ancient Chinese characters. Due to drilling and burning before and during divination, and the long-term underground corrosion, as well as excavation, transportation, and circulation after their excavation, about 90% of the OBs have been fragmented and are now scattered in different collections around the world [1].
As precious cultural relics, many OBs were circulated among various collectors and antique dealers in the initial period after their discovery in 1899. Limited by communication and dissemination methods at that time, the same OBs might have been repeatedly published in different publications at different times in different countries/regions, which led to the phenomenon of OB duplicates, referring to the fragments being repeatedly published. Some Oracle Bones further fragmented during circulation; on the other hand, as OBI research advances, some fragmentary OBs might have been rejoined by OBI domain experts and republished again. As such, OB duplicates exhibit both one-to-one and one-to-many image matching relationships. Although domain experts have manually found many duplicates in their research and compilation process, considering the huge cardinality of OB fragments (more than 160,000), AI-enabled OB duplicate discovery becomes imperative.
OBI is carved writing, its main publication format is rubbing, in which people place a paper onto the surface of the bones, then use Rubbing (with inks) to copy the carved inscriptions. Domain experts can also reproduce (copy) the carved inscriptions by hand, which is named Manual Oracle Bone Inscriptions Copy. Comparing the two formats, manual OBI copies rely on domain knowledge but has no background noises, whereas OBI rubbings often contains obvious noise disturbance, although domain knowledge is not required. Both formats can keep the original sizes of the Oracle Bones, which is not possible when using cameras. Until today, OBI rubbings and manual copies remain the main materials for OBI research. In 2022, the largest collection of manual OBI copies was published [2], for which a large team of OBI researchers invested 10 years to create high-quality manual OBI copies for around 60,000 OBs.
With this new collection at hand, in this work we aim to devise a comprehensive framework for discovering OB duplicates at large-scale. Since different domain experts have slightly different copying styles for the same OBIs (such as variations in pen movement, stroke thickness, brush pressure), finding OB duplicates can be essentially formulated as a content-based image retrieval (CBIR) [3, 4, 5] or image matching [6, 7, 8] task.
Contributions. To our knowledge, this work is among the first technical efforts that investigate AI-enabled Oracle Bone duplicates discovery. We design OBD-Finder, an explainable coarse-to-fine Oracle Bone duplicates discovery framework that successively utilizes unsupervised low-level key feature points matching and high-level content (character) similarity for ranking the OB duplicate candidates.
We conduct extensive experiments on a large dataset of OB copies, and compare our framework with state-of-the-art CBIR and image matching methods. We demonstrate that our approach achieves Top-K recall performance comparable to state-of-the-art methods, but with significantly accelerated computational efficiency and substantially reduced GPU memory consumption. Additionally, it attains the highest simplified mean reciprocal rank (MRR) scores for both Top-5 and Top-15 retrieval results, demonstrating that our approach excels at prioritizing correct matches. We have deployed our approach in real-world applications, where we have successfully identified over 60 pairs of new Oracle Bone duplicates, which have been missed by domain experts for decades.
2 Related Work
AI-enabled ancient manuscript fragments retrieval and rejoining have very important real-world applications [9, 1, 10]. Our domain-specific Oracle Bone (OB) duplicates discovery problem can be essentially formulated as a image retrieval or image matching task. In the following, we provide a brief overview of the content-based image retrieval (CBIR) and image matching techniques assessed in this study.
CBIR techniques.
The Smooth-AP loss [3] provides a differentiable approximation for Average Precision (AP) to enable end-to-end training for ranking-based CBIR tasks. HashNet [11] adopts a sign activation function for binarizing the K-dimensional deep feature representation into K-bit binary hash code for retrieval tasks. In this paper, we use the latest transformer-based implementation of HashNet [12]. HybridHash [5] is a hybrid deep hashing architecture combining self-attention and convolutional layers for enhancing image retrieval performance. We note that, since there also exists one-to-many mapping relationship in OB duplicates, contour/shape-based retrieval methods are not fully applicable in our task.
Image matching techniques.
Image matching, also known as feature/keypoint matching, aims to establish correspondences between points in two images depicting the same scene or object. SIFT (Scale-Invariant Feature Transform) [13] is a classic algorithm for image matching. In recent years, deep learning based techniques have substantially improved the state-of-the-art in this field. LoFTR [6] utilizes Transformer self-attention and cross-attention mechanisms to establish coarse-level matches of the keypoints, followed by fine-level match on the cropped local windows for each coarse match. OmniGlue [7] leverages vision foundation models for boosting generalization to unseen domains. It also uses self- and cross-attention for establishing intra- and inter-image connectivity graphs to enhance correspondence estimation. MINIMA [8] is a latest image matching technique that contains a simple data engine for freely generating multiple modalities for images to pre-train modality invariant image matching model. In this work, we compare with the best pre-trained MINIMA model.
Despite their technical advancement and substantial performance improvement, OmniGlue essentially depends on foundation models for accurate image matching, while MINIMA relies on “heavy” pre-training on generated multi-modal data, and both of them have significantly slower inference speed, compared to our proposed framework. Moreover, all the above image matching methods depend on low-level feature matching, but lack high-level semantic guidance. In comparison, in this work we seamlessly integrate unsupervised low-level feature matching and high-level character-centric semantic-aware content matching, offering outstanding retrieval and matching performance, fast inference speed, and strong interpretability for OB duplicates discovery.
3 Methodology
Framework.
As can be seen from Figure 1, we propose a progressive coarse-to-fine Oracle Bone duplicate discovery framework, namely OBD-Finder, which combines unsupervised low-level keypoint matching with high-level, character-centric content-based image matching. Keypoint matching operates at low-level visual feature scale, which can prune out candidates with low degree of match in the initial stage, but lacks explicit semantic supervision and interpretability. Our framework bridges this gap by first grouping the keypoints based on their association with the character regions, then assesses the global matching degree between the two groups of keypoints via character-level visual content similarity computation. This dual matching mechanism enhances Oracle Bone duplicates discovery accuracy through a progressive coarse-to-fine refinement manner, by effectively and seamlessly integrating both low-level keypoint and high-level character-based semantic cues, resulting in more accurate and semantic-aware image matching. Our framework consists of four subsequent steps:
Figure 1: The overall framework of OBD-Finder for Oracle Bone duplicates discovery.
  1. (1)
    Feature Extraction. We perform unsupervised feature points/keypoints extraction on the OBs using a pre-trained model [14].
  2. (2)
    Feature Matching. We next apply unsupervised keypoints mapping between the two OB images using a pre-trained model [15]. Candidate with low overall matching degrees will be filtered out in this step.
  3. (3)
    Coordinate Alignment. After obtaining the correspondence between the keypoints in feature matching, we apply affine transformations for each image pair, in which we map the coordinates of the image with fewer feature points to the other image.
  4. (4)
    Character-level Content Similarity. we first localize the Oracle Bone characters in each image, using a text detector [16]. Given that the coordinate systems of two images are aligned, for each character in the smaller image, we search for the overlapped characters in the counterpart image, then compute the content similarity between the overlapped characters, using a simple Siamese network model [17].
Key Features.
(i) OBD-Finder is a progressive coarse-to-fine framework that seamlessly proceeds from low-level keypoints matching to high-level semantic-aware content similarity computation, resulting in very accurate OB duplicate discovery. (ii) It is a transparent pipeline with strong interpretability. (iii) It is unsupervised and almost training-free, only requiring little annotation effort. (iv) It is highly efficient, compared to state-of-the-art image matching methods, which we will demonstrate in the empirical studies.
4 Experiments
Dataset.
We use the 60,000 manual Oracle Bone copies from the Oracle Bone Inscription Copy Series [2] as the retrieval database, which have been divided into 25 OB groups by the OBI community based on dating and writing style. Two OBI domain experts who are co-authors of this work (Prof. Yi Men and Ms. Yingqi Chen) have manually collected 150 pairs of OB duplicates published in the OBI literature, which belong to different OB groups and will be used the query images. For each pair of images, we perform OB duplicate checking in the corresponding OBI group. For OBI character localization using EAST [16] and character-level content matching using Siamese network [17], OBI domain experts have manually annotated the character regions and categorization for 500 OBs.
Evaluation Metrics.
We use two evaluation metrics: Recall@K and a simplified mean reciprocal rank measure MRR@K. Recall@K measures how many relevant items were successfully retrieved in the Top-K results, but doesn’t consider their rankings. To address this limitation, we also adopt a modified version of MRR@K, since our dataset only contains one correct match per query, we simply average the rank position of the correct answer in the Top-K results across all query images, abbreviated as Rank@K.
Table 1: Comparing OBD-Finder with CBIR methods (Recall@K performance).
Method Recall@1 Recall@5 Recall@10 Recall@15 Recall@20
Smooth-AP [3] 34.1 62.9 73.8 78.8 82.8
Proxy-Anchor [4] 71.5 77.0 84.1 86.6 91.9
HashNet [12] 57.6 63.8 69.3 78.4 85.8
HybridHash [5] 60.8 66.8 73.2 79.3 88.3
Ours (OBD-Finder) 80.0 85.3 90.4 94.3 98.0
Results.
We first compare our approach with content-based image retrieval (CBIR) methods. As reported in Table 1, our approach (OBD-Finder) consistently obtains the best recall performance than the CBIR methods.
We next compare our approach (OBD-Finder) with state-of-the-art image matching techniques. As shown in Table 2, it exhibits recall performance on par with the latest image matching methods. In Figure 2, we showcase image matching results of different methods. Moreover, as reported in Tables 3, our approach obtains the best Rank@K scores in the Top-5 and Top-15 retrieval results, demonstrating that it excels at prioritizing correct matches.
Table 2: Comparing OBD-Finder with image matching methods (Recall@K).
Method Recall@1 Recall@5 Recall@10 Recall@15 Recall@20 Recall@25
SIFT [13] 33.3 40.0 46.8 53.3 66.6 73.2
LoFTR [6] 73.6 75.4 81.3 86.0 92.3 98.3
OmniGlue [7] 82.5 86.0 91.2 95.6 98.2 100
MINIMA [8] 84.4 90.5 94.7 98.2 100 100
Ours (OBD-Finder) 80 85.3 90.4 94.3 98 100
Figure 2: Case studies for image matching results of different methods.
Table 3: Comparing OBD-Finder with image matching methods (Rank@K).
Method Rank@5 Rank@10 Rank@15 Rank@20 Rank@25
SIFT [13] 1.6 2.70 4.16 6.15 7.74
LoFTR [6] 1.09 1.88 2.74 3.76 5.22
OmniGlue [7] 1.13 1.51 2.05 2.52 2.89
MINIMA [8] 1.36 1.66 2.07 2.38 2.38
Ours (OBD-Finder) 1.06 1.63 2.00 2.61 2.93
Table 4: Comparisons on Inference speed, FPS, and GPU memory consumption.
Method Recall@20 Inf. Speed (s/pair) FPS (pair/s) GPU (MiB)
SIFT [13] 66.6 0.017 59 N/A
LoFTR [6] 92.3 3.6 0.28 3502
OmniGlue [7] 98.2 45 0.02 23612
MINIMA [8] 100 1 1 14890
Ours (OBD-Finder) 98 0.021 47.62 5215
Finally, in Tables 4, we can observe OBD-Finder has significantly faster inference speed than state-of-the-art image matching algorithms (about 40 times faster), but substantially less GPU consumption (about 1/3 of their GPU usages). Therefore, OBD-Finder is accurate and efficient at prioritizing correct matches.
Real-world Deployment.
Besides empirical study, we also perform OB duplicate checking for other OBs in each category, where we have successfully discovered over 60 groups of new OB duplicates, which have been verified by OBI domain experts (Prof. Yi Men and Ms. Yingqi Chen, who are also co-authors of this work). In the project homepage, we showcase samples of the new OB duplicates that we have discovered in real-world deployment.
5 Conclusion
For AI-enabled Oracle Bone duplicates discovery, we integrate unsupervised low-level feature matching with high-level character-based visual content matching for accurately and efficiently identifying the correct OB duplicates. We have discovered over 60 pairs of new OB duplicates in real-world deployment. In future work, we will jointly use the dual modalities of rubbing and manual copies to conduct multi-modal OB duplicates discovery.